loralei:

The table shows the weights of giant pumpkins grown in two towns, Cedartown and Baytown. Match each phrase to its value. Weight of Cedartown Pumpkins (pounds) Weight of Baytown Pumpkins (pounds) 611 584 753 622 575 713.5 685.5 579 718 644 20.2 40 40.2 60.4 575 628.5 668.5 713.5 the average weight of pumpkins grown in Cedartown arrowRight the average weight of pumpkins grown in Baytown arrowRight the mean absolute deviation of the pumpkin weights in Cedartown arrowRight the mean absolute deviation of the pumpkin weights in Baytown arrowRight

Extrinix:

Can you provide a screenshot of the chart for us please.

HaydenwH:

with the numbers being all in a row, we can not direct you into working it, if the numbers arent seperated.

HaydenwH:

do you know how to find average?

HaydenwH:

hen finding an average of set numbers, you would begin by adding all of the numbers in a set, and then dividing the sum by the amount of numbers in the set

Extrinix:

So basically take all of the numbers of Cedartown and add them, then divide the sum of those numbers by 5 \(611+753+575+685.5+718=X\) so \(3342.5=X\) then we take that and divide it by 5 so \(\dfrac{3342.5}{5}\) that gives you youre answer for the first question. You can do the same to find question 2 that way. now to find the mean absolute deviation of the two towns, do this: I made up an equation for this, \(x= \dfrac{v-m}{t}\) now to do mean absolute deviation, you find the mean listed in the first 2 questions, then take that and find how far from the mean each number is, so for instance the first number listed in Cedartown, \(611 \rightarrow \dfrac{611-668.5}{5}\) now, do that for each value then add all of your answers together and divide by 5 \(\dfrac{477.3+v2+v3+v4+v5}{5}=x\) once you solve the last part, you have your answer for Cedartown, then do the same for Baytown.

Extrinix:

If you need any help, just leave a comment and I'll assist you.